OpenStudy (anonymous):

solve the differential equation: \[\frac{dx}{dt}=k(x-a)e^{-\frac {(x-a)^2}{b}}.\]

OpenStudy (anonymous):

anyone ??

OpenStudy (anonymous):

wt do we have to do?

OpenStudy (dumbcow):

that looks a lot like the normal distribution function

OpenStudy (anonymous):

yes exactly @dumbcow....@doncatch you have to solve the differential equation and express x interms of t

OpenStudy (mr.math):

This is a separable DE. \[\frac{dx}{dt}=k(x-a)e^{-\frac {(x-a)^2}{b}} \implies \frac{dx}{k(x-a)}e^{\frac {(x-a)^2}{b}}=dt.\] Integrate both sides now.

OpenStudy (mr.math):

You probably know that the integral on the LHS can't be found in terms of elementary functions. So you can write \[\large \int_0^x \frac{e^{\frac {(\lambda-a)^2}{b}}}{{k(\lambda-a)}} d\lambda=t+c.\]